OpenStudy (anonymous):

find the second derivative of y^2=x^2+2x

OpenStudy (amistre64):

just derive it like you normally would, implicits aint a new thing :)

OpenStudy (anonymous):

\[1/y^3\] is this the answer

OpenStudy (amistre64):

i doubt it

OpenStudy (amistre64):

what steps did you take to derive it?

OpenStudy (anonymous):

differentiating both sides wrt x then double derivative

OpenStudy (amistre64):

ok, what is the first derivative wrtx of: y^2=x^2+2x

OpenStudy (anonymous):

(x+1)/y

OpenStudy (amistre64):

2y y' = 2x + 2 y' = (x+1)/y ; yep, i agree and the derivative of that is?

OpenStudy (anonymous):

\[(y-(x+1)dy/dx)/y ^{2}\]

OpenStudy (amistre64):

y'' = [y(x+1)' - y'(x+1)]/y^2 y'' = (y - xy'+y')/y^2 y'' = (y^2/y - x(x+1)/y +(x+1)/y)/y^2 y'' = (y^2 - x(x+1) +(x+1))/y^3 y'' = (y^2 - x^2-x+x+1)/y^3 y'' = (y^2 - x^2 +1)/y^3 is what I get

OpenStudy (amistre64):

forgot to drag the negative thru in the second step tho

OpenStudy (amistre64):

y'' = [y(x+1)' - y'(x+1)]/y^2 y'' = (y - xy'-y')/y^2 y'' = (y^2/y - x(x+1)/y -(x+1)/y)/y^2 y'' = (y^2 - x(x+1) -(x+1))/y^3 y'' = (y^2 - x^2-x-x-1)/y^3 y'' = (y^2 - x^2-2x-1)/y^3 looks a bit better maybe

OpenStudy (anonymous):

can't we put value of y^2 from the question? if so the ans is 1/y^3